# Learning Part-aware 3D Representations by Fusing 2D Gaussians and Superquadrics

### v1

@@ III-B Optimizing the Hybrid Representation
In this section, we introduce how to optimize the hybrid representation. By minimizing the rendering loss across multi-view images only, we found it unstable to position the hybrid blocks (primitives) accurately. Thus, we introduce several regularizations and perform two optimization stages (block-level and point-level optimizations).
III-B1 Block-level Optimization
@@
Besides the image rendering loss, we employ regularizers consisting of coverage, overlap, parsimony, and opacity entropy. The coverage loss encourages hybrid blocks to cover only the meaningful regions; the overlap loss prevents different blocks from overlapping; the parsimony loss regularizes the number of existing blocks; the opacity entropy encourages opacities of hybrid blocks close to binary.
Rendering Loss. Based on Eq. 7, we can obtain the rendered image $\hat{\mathbf{I}}$ associated with a given camera pose. We calculate the rendering loss between $\hat{\mathbf{I}}$ and the ground truth image $\mathbf{I}$ by associating $L_{1}$ with a D-SSIM term[18]:
$L_{\text{ren}}=(1-\lambda)\ L_{1}+\lambda L_{\text{D-SSIM}}.$ (8)
@@
Coverage Loss. This loss ensures that the block set $\{\mathcal{B}_{1}\dots\mathcal{B}_{M}\}$ covers the object while preventing it from extending outside the object. To determine the occupancy relationship in 3D space, we first define the approximate signed distance of a 3D point $x$ to the $i$-th block, $D_{i}(x)=1-\Psi_{i}(x)$. Here, $\Psi_{i}$ denotes the superquadric inside-outside function [6] associated to the block $i$. Consequently, $D_{i}(x)\geq 0$ if the point $x$ lies inside the $i$-th block, and $D_{i}(x)<0$ if $x$ is outside of the block. Similar to NeRF [14], we sample a ray set $\mathcal{R}$ using ray-casting based on camera poses. Given the object mask, we associate each ray $r\in\mathcal{R}$ with a binary label, such that $l_{r}=1$ if the ray $r$ is inside the mask, otherwise $l_{r}=0$. The coverage loss is defined as:
@@
Here, $L_{\text{cross}}$ encourages rays to intersect with blocks , while $L_{\text{non-cross}}$ discourages rays from intersecting with blocks. Intuitively, this means that there is at least one in sampled points ${\mathcal{X}}_{r}$ along the ray $r$ inside the mask that lies within a certain block, while all points along the ray outside the mask should not lie within any block. This process can be expressed as:
@@
where $x_{j}^{r}$ means the $j$-th sampled points along the ray $r$ and =$\mathcal{M}=\{1,\dots M\}$.
Overlap Loss. We add a regularization term to encourage individual blocks to be non-overlapping. Considering the difficulty in directly calculating the overlap of locks, we use a Monte Carlo method similar to  [2]. Practically, we sample multiple 3D points in space and penalize each point for being inside more than $k$ blocks. Based on the definition of the superquadric inside-outside function, we devise a soft occupancy function:
@@
where $\gamma$ is a temperature parameter. Therefore, given a set of $N$ sampled 3D points $\Omega$, the overlap loss can be expressed as:
$L_{\text{over}}=\frac{1}{N}\sum_{x\in\Omega}\text{ReLU}\left(\sum_{i\in\mathcal{M}}\mathcal{O}_{i}^{x}-k\right).$ (13)
Parsimony Loss. To promote the use of the minimal number of blocks and achieve parsimony in decomposition, we add a loss term that regularizes the number of active blocks. This loss is defined as:
$L_{\text{par}}=\frac{1}{M}\sum_{i\in\mathcal{M}}\sqrt{\tau_{i}}.$ (14)
Opacity Entropy Loss. Naturally, during optimization, the opacity of blocks inside the object region tends to approach 1, while the opacity of blocks outside the object region tends to be 0. To facilitate this, we associate the opacities of the blocks with the labeled rays. Specifically, we employ a cross-entropy loss, $L_{ce}$, between the opacity and mask labels as follows:
${\mathcal{L}}_{\text{opa}}({\mathcal{R}})$ $=\frac{1}{|{\mathcal{R}}|}\sum_{r\in{\mathcal{R}}}L_{ce}\left(\max_{i\in\mathcal{M}}\tau_{i}(x^{r}),l_{r}\right)~{},$ (15)
$\left\{x^{r}\in{\mathcal{X}}_{r}\mid D_{i}(x^{r})\leq 0\right\}.$
Thus, the total loss is the weighted summation of the above terms:
$L=L_{\text{ren}}+\lambda_{\text{cov}}L_{\text{cov}}+\lambda_{\text{over}}L_{\text{over}}+\lambda_{\text{par}}L_{\text{par}}+\lambda_{\text{opa}}L_{\text{opa}}.$ (16)
Adaptive Number of Blocks. Given that the number of parts varies by scene, we allow the number of blocks to adjust during optimization adaptively. When the opacity of an existing block falls below a threshold $t$, this block is immediately removed. Additionally, we explore a block-adding mechanism using point clouds that are randomly generated within the scene area. Specifically, we employ DBSCAN [54] to cluster point clouds within the perspective of the object masks and not covered by any blocks. A new primitive is then introduced at the center of each point cloud cluster. The other parameters of these new blocks are initialized randomly.
@@
III-B2 Point-level Optimization
The primitive-based hybrid representation effectively decomposes the shape into parts but underperforms with complex objects. To address this, we perform a refinement stage for precise geometry fitting. In this stage, we free the constraint that Gaussian ellipses must be at the face of the superquadric mesh, allowing Gaussians to move freely to fill in the complex regions. Moreover, to prevent Gaussian ellipses from one block passing through another, we impose a new constraint to minimize the negative signed distance of each Gaussian entering other blocks:
$L_{\text{enter}}=\frac{1}{N}\sum_{x\in\Omega}\sum_{m\in\mathcal{M}\setminus\{\delta\}}\text{ReLU}(-D_{m}(x)),$ (17)
where $\Omega$ is the sampled Gaussian set, and $\delta$ represents the identifier of the block to which the Gaussian $x$ belongs. After this stage, as shown in Fig. 2, the reconstruction is more precise and better fits the original shape.
Figure 3: Qualitative comparisons on DTU [55]. We compare our approach with state-of-the-art baselines on the DTU dataset with the background removed. The last two columns show our block-level and point-level reconstructions, respectively. Our method is the only one that provides reasonable 3D part decomposition while capturing detailed geometry.
Figure 4: Qualitative comparisons on ShapeNet [55]. We compared our approach with state-of-the-art baselines across four categories. The last two columns display our block-level and point-level reconstructions, respectively. Our method uniquely provides reasonable 3D part decomposition while simultaneously capturing detailed geometry.
Figure 5: Qualitative results on BlenderMVS [56] and self-capatured data. We demonstrate the RGB renderings and decomposed parts from novel views across a variety of objects. The first five examples are from the BlenderMVS dataset, and the remaining examples are from our own captured scenes.
Figure 6: Ablation studies on key strategies. The block-level visual comparisons illustrate the impact of adopting our proposed strategy. The first row indicates results without the strategy, while the second row shows results with the strategy implemented.
TABLE I: Quantitative comparison on DTU [55]. The Chamfer Distance (CD) between the 3D reconstruction and the ground-truth is reported in 15 scenes. The best results are bolded, and the average numbers of primitives found (#P) that are greater than 10 are underlined.
Chamfer Distance (CD) per scene Mean Mean
Method Input S24 S37 S40 S55 S63 S65 S69 S83 S97 S105 S106 S110 S114 S118 S122 CD #P
EMS [1] GT point cloud 6.32 3.54 2.99 4.30 4.16 4.01 3.75 3.24 4.97 4.34 4.16 7.62 7.58 4.46 4.03 4.65 7.7
MBF [41] GT point cloud 3.12 2.66 3.84 2.54 1.59 2.11 2.19 2.01 2.32 2.45 2.17 2.12 3.83 2.02 2.55 2.50 34.1
@@ III-B2 Point-level Optimization
TABLE II: Quantitative comparisons on DTU [55]. We compare our method to PartNeRF, DBW, and 2DGS in terms of PSNR, SSIM, LIPPS, Chamfer Distance, and reconstruction time. All running time measurements were conducted on a single GTX RTX3090 GPU. Excluding the results of 2DGS, the best results are in bold font, and the second-best results are underlined.
Method Structure CD $\downarrow$ PSNR $\uparrow$ SSIM  $\uparrow$ LIPPS $\downarrow$ Time $\downarrow$
2DGS [17] 0.96 34.07 0.99 0.019 $\sim$ 18 mins
PartNeRF [15] 9.59 17.97 0.77 0.246 $\sim$ 8 hours
DBW [2] 4.73 16.44 0.75 0.201 $\sim$ 2 hours
Ours (Block-level) 4.19 19.84 0.82 0.189 $\sim$ 30 mins
Ours (Point-level) 1.05 35.04 0.99 0.015 $\sim$ 50 mins
TABLE III: Quantitative comparison on ShapeNet [57]. The Chamfer Distance (CD) between the 3D reconstruction and the ground-truth is reported on objects of four categories. The best results are bolded, and the second-best results are underlined.
@@
TABLE IV: Quantitative comparison on ShapeNet [57]. We report the PSNR, SSIM, LIPPS, Chamfer Distance, and reconstruction time for our method, PartNeRF, as well as DBW and 2DGS across four categories. Excluding the results of 2dgs, the best results are bolded, and the second-best results are underlined.
@@
Method Structure Airplane Table Chair Gun Airplane Table Chair Gun Airplane Table Chair Gun Airplane Table Chair Gun
@@
Ours (Block-level) 2.47 2.15 2.32 1.78 26.93 26.92 23.91 27.97 0.936 0.911 0.891 0.942 0.082 0.139 0.126 0.072
@@
IV Experiments
IV-A Experimental Settings
Datasets. We conduct evaluation experiments on two public datasets: DTU [55] and ShapeNet [57]. DTU is an MVS dataset comprising 80 forward-facing scenes, with 49-60 images captured for each scene. For evaluation, we utilize 15 scenes used in 2DGS [17]. We downsample the images to a resolution of 400 $\times$ 300 for efficiency. To validate the effectiveness of our method in segmenting parts of man-made objects, we create a ShapeNet dataset that includes four categories: Chair, Table, Gun, and Airplane. Each category comprises 15 distinct objects, offering diverse instances for comprehensive evaluation. We randomly sample 100 camera poses on the sphere centered on each shape and render images at a resolution of 400$\times$400. The rendered images are equally split into training and testing sets. Additionally, to explore our method’s potential in handling real-life data, we present qualitative results on the BlenderMVS dataset [56] and self-captured scenes. For BlenderMVS data, we utilize standard camera poses. For self-captured scenes, we recover the camera poses using COLMAP [21] and normalize them following IDR [58]. All images are resized to 400 $\times$ 300 pixels. Object masks are obtained using Grounded SAM [59].
Baselines. We compare our method with four state-of-the-art works on 3D shape decomposition: EMS [1], MonteboxFinder (MBF) [41], PartNeRF [15] and DBW [2]. MBF and EMS are applied to point clouds, utilizing cuboids and superquadrics, respectively, to fit the 3D points. The input point clouds are sampled from either the original 3D ground truth (GT) or the mesh reconstructed by the state-of-the-art MVS method Neus [27]. We adhere to the testing procedure outlined by DBW [2] and randomly sample 5K and 200K points from the GT meshes as the input for EMS and MBF, respectively. Similar to our problem setting, PartNeRF and DBW use images directly as input to construct structural 3D representations. To ensure a fair comparison, we discarded the pre-defined ground plane in DBW since the test scenario consists solely of foreground objects. For PartNeRF, we perform instance-specific training and set the same number of parts as in our method. Additionally, we compare our method with 2DGS [17], the state-of-the-art Gaussian Splatting reconstruction method.
Metrics. We evaluate from 3D reconstruction quality, view synthesis, shape parsimony, and reconstruction time. The 3D reconstruction quality is measured by the Chamfer Distance (CD) between the recovered geometry and GT, reflecting the accuracy of 3D reconstructions. View synthesis uses three standard metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [60], and Learned Perceptual Image Patch Similarity (LPIPS) [61]. Shape parsimony and reconstruction time are measured by the number of parts and running time, where fewer parts facilitate easier part editing or manipulation.
IV-B Implementation Details
We improve the 2DGS [17] architecture by adding a non-learnable part attribute to each Gaussian component to enforce the constraint in Eq. 17 during refinement optimization. We employ regularization in 2DGS to achieve better geometric reconstruction, including depth distortion maps, depth maps, and normal maps. Additionally, we introduce a mask cross-entropy to filter out extra noise Gaussians. To extract the meshes from 2D Gaussians, we use truncated signed distance fusion (TSDF) to fuse the rendered depth map, utilizing Open3D [62]. In the ShapeNet benchmark, we observed a common degradation in the mesh reconstruction by 2DGS, where the meshes of objects contain many holes. This issue arises due to the lack of textures in ShapeNet images, causing 2DGS to use large-scale Gaussians. This leads to depth inconsistencies from multiview, resulting in holes in the constructed mesh. To address this, we apply scale regularization to limit the scale of large Gaussian components. We validate this design in Sec. IV-D.
@@ IV-B Implementation Details
The same hyperparameters are used for all experiments. We set the initial number of primitives $M$ to 8. In the hybrid representation, each superquadric mesh is a level-2 icosphere (320 triangular face). Each triangular face contains 100 Gaussians, with a scale parameter $c$ of 0.15 and a block filter threshold $t$ of 0.2. The temperature parameter $\gamma$ in Eq. 12 is 0.005 and the overlapping number $k$ in Eq. 13 is 1.95. We train the hybrid representation for 30k iterations, followed by refinement optimization for another 30k iterations. During block-level optimization, the block-adding operation is carried out at the 5000th and 10000th iterations. The point clouds used for block-adding are randomly initialized within the space $[-1,1]^{3}$. The number of sampled points in each ray is 2048 in Eq. 9 and Eq. 15. The loss weights $\lambda_{\text{cov}}$, $\lambda_{\text{over}}$, $\lambda_{\text{par}}$, and $\lambda_{\text{opa}}$ are set to 10, 1, 0.01, and 0.01, respectively.
IV-C Evaluations
IV-C1 DTU Benchmark
We utilize the official DTU Chamfer Distance (CD) evaluation [55] to compare the performance between our method and these state-of-the-art 3D decomposition techniques, as shown in Tab. I. Additionally, we present the input type and the average number of primitives used. Our method consistently outperforms in all scenes compared to prior works. Even without refinement optimization, our method still provides satisfactory CD errors. PartNeRF has the largest CD errors in the DTU dataset. We speculate that this is due to PartNeRF’s difficulty in handling forward scenes. In contrast, MBF achieves a low CD error but at the cost of using significantly more primitives, leading to over-decomposition and the inclusion of meaningless parts. Although EMS and DBW utilize a reasonable number of primitives, they yield higher CD errors than the proposed method.
We compare the quality of view synthesis and reconstruction time with PartNeRF, DBW, and 2DGS in Table II. Even without point-level optimization, our method achieves superior image rendering quality compared to other state-of-the-art methods. With the addition of refinement, both image rendering quality and reconstruction accuracy are further improved, nearing the performance of the state-of-the-art Gaussian-based reconstruction method, 2DGS [17]. However, 2DGS can only produce unstructured meshes. Notably, our model demonstrates excellent efficiency, with a reconstruction speed approximately 10X faster than the implicit reconstruction method PartNeRF and over 2X faster than DBW.
Figure 3 presents visual comparisons of our method with existing state-of-the-art baselines. Even with ground-truth 3D points, EMS and MBF still struggle to extract reasonable compositions. PartNeRF reconstructs reasonable geometry but fails to decompose the scenes meaningful. DBW experiences frequent failures, particularly when parts of the geometry are not reconstructed due to the absence of primitives in these regions at initialization. In contrast, our approach consistently produces interpretable 3D decompositions (second-to-last column) and, through refinement, achieves more detailed structures (last column).
IV-C2 ShapeNet Benchmark
We evaluate our approach on four categories of ShapeNet [57]. Since MBF [41] fails on nearly all objects of the ShapeNet benchmark, we compare only against EMS [1], PartNeRF [15] and DBW [2]. Fig. 4 presents the qualitative comparisons of the 3D decomposition of eight object shapes, with two instances per category. Tab. III and Tab. IV deliver detailed quantitative comparisons on reconstruction quality, novel view synthesis, and primitives parsimony. For clarity, the chamfer distance is multiplied 100 times in the Shapenet benchmark.
EMS yields large Chamfer distances, especially in the chair and table categories, despite fitting superquadrics from GT 3D points. Similarly, DBW often fails to recover complete geometry in these categories, resulting in significant reconstruction errors. However, it performs better with guns and airplanes due to their simpler shapes. The non-convex nature of chairs and tables poses substantial challenges for EMS and DBW in generating reasonable geometry. PartNeRF benefits from NeRF’s geometric reconstruction capabilities to produce accurate geometry and good novel view renderings. However, it has two main disadvantages. On the one hand, PartNeRF cannot adaptively adjust the primitives, resulting in a fixed number of parts. Estimating the number of parts in advance is impractical. On the other hand, its use of ellipsoidal primitives limits its ability to decompose meaningful parts, such as separating the chair seat and tabletop into several components, as shown in the three columns of Fig. 4.
Conversely, the superior performance of our approach is demonstrated across all categories, visually and quantitatively. In addition, thanks to Gaussian’s powerful rendering capabilities, clear advantages in novel view synthesis are observed over PartNeRF and DBW, as shown in Tab. IV. The quantitative comparisons against 2DGS [17] further validate the capability of our method on reconstruction accuracy. Our method outperforms 2DGS in the chair and airplane categories and achieves comparable performance in the other two categories.
TABLE V: Ablation study on the ShapeNet [57] chair benchmark. We report Chamfer Distance (CD), rendering metrics and the number of primitives (#P) of block-level reconstruction and point-level reconstruction. The best results are in bold font, and the second-best results are underlined.

### v2

@@ 3.2 Optimizing the Hybrid Representation
This section describes the optimization of the hybrid representation. We found that minimizing the rendering loss across multi-view images alone led to instability in positioning hybrid blocks (primitives) accurately. Therefore, we introduce several regularizations and perform two optimization stages: block-level and point-level optimization.
3.2.1 Block-level Optimization
@@
In addition to the image rendering loss, we employ regularizers consisting of coverage, overlap, parsimony, and opacity entropy. The coverage loss encourages hybrid blocks to cover only meaningful regions; the overlap loss prevents blocks from overlapping; the parsimony loss regularizes the number of existing blocks; and the opacity entropy encourages block opacities close to binary.
Rendering Loss. Based on the rendering process of Gaussian spaltting [18], the rendering loss is calculated by associating $L_{1}$ with a D-SSIM term[18] :
${\mathcal{L}}_{\text{ren}}=(1-\lambda)\ L_{1}+\lambda L_{\text{D-SSIM}}.$ (5)
@@
Coverage Loss. This loss ensures that the block set $\{\mathcal{B}_{1}\dots\mathcal{B}_{M}\}$ covers the object while preventing it from extending beyond the object’s boundaries. To determine the 3D occupancy field based on the blocks, we first define the approximate signed distance of a 3D point $x$ to the $i$-th block, $D_{i}(x)=1-\Psi_{i}(x)$. Here, $\Psi_{i}$ denotes the superquadric inside-outside function [2] associated with the block $i$. Consequently, $D_{i}(x)\geq 0$ if the point $x$ lies inside the $i$-th block, and $D_{i}(x)<0$ if $x$ is outside the block. Inspired by NeRF [25], we sample a ray set $\mathcal{R}$ using ray-casting based on camera poses. Given the object mask, we associate each ray $r\in\mathcal{R}$ with a binary label: $l_{r}=1$ if the ray $r$ is inside mask, otherwise $l_{r}=0$. The coverage loss is defined as:
@@
Here, $L_{\text{cross}}$ encourages rays to intersect with blocks, while $L_{\text{non-cross}}$ discourages rays from intersecting with blocks. They are defined as:
@@
where $x_{j}^{r}$ means the $j$-th sampled points along the ray $r$ and =$\mathcal{M}=\{1,\dots M\}$. Intuitively, this means that there is at least one in sampled points ${\mathcal{X}}_{r}$ along the ray $r$ inside the mask that lies within a certain block, while all points along the ray outside the mask should not lie within any block.
Overlap Loss. We add a regularization term to encourage individual blocks to remain non-overlapping. Given the difficulty of directly calculating block overlap, we use a Monte Carlo method similar to  [27]. Practically, we sample multiple 3D points in space and penalize each point that lies inside more than $k$ blocks. Based on the definition of the superquadric inside-outside function, we devise a soft occupancy function:
@@
where $\gamma$ is a temperature parameter. Therefore, given a set of $N$ sampled 3D points $\Omega$, the overlap loss is expressed as:
${\mathcal{L}}_{\text{over}}=\frac{1}{N}\sum_{x\in\Omega}\text{ReLU}\left(\sum_{i\in\mathcal{M}}\mathcal{O}_{i}^{x}-k\right).$ (10)
Parsimony Loss. To promote the use of the minimal number of blocks and achieve parsimony in decomposition, we add a loss term that regularizes the block opacity ($\tau$). This loss is defined as:
${\mathcal{L}}_{\text{par}}=\frac{1}{M}\sum_{i\in\mathcal{M}}\sqrt{\tau_{i}}.$ (11)
Opacity Entropy Loss. Naturally, during optimization, the opacity of the block inside the object region tends to approach 1, while the opacity of the block outside the object region tends to approach 0. To facilitate this, we associate the opacities of the blocks with the labeled rays. Specifically, we use a cross-entropy loss $L_{ce}$, between the opacity and mask labels as follows:
${\mathcal{L}}_{\text{opa}}({\mathcal{R}})=\frac{1}{|{\mathcal{R}}|}\sum_{r\in{\mathcal{R}}}L_{ce}\left(\max_{i\in\mathcal{M}}\tau_{i}(x^{r}),l_{r}\right)~{}.$ (12)
Here, only points $x^{r}$ inside the blocks are sampled, e.g., $\left\{x^{r}\in{\mathcal{X}}_{r}\mid D_{i}(x^{r})\leq 0\right\}$.
The total loss is the weighted summation of the loss terms described above:
${\mathcal{L}}={\mathcal{L}}_{\text{ren}}+\lambda_{\text{cov}}{\mathcal{L}}_{\text{cov}}+\lambda_{\text{over}}{\mathcal{L}}_{\text{over}}+\lambda_{\text{par}}{\mathcal{L}}_{\text{par}}+\lambda_{\text{opa}}{\mathcal{L}}_{\text{opa}}.$ (13)
Adaptive Number of Blocks. Given that the number of parts varies across scenes, we allow the number of blocks to adjust during optimization adaptively. When the opacity of an existing block falls below a threshold $t$, this block is immediately removed. Additionally, we explore a block-adding mechanism using point clouds that are randomly generated within the scene area. Specifically, we use DBSCAN [34] to cluster initial point clouds (e.g., accessible by-products of COLMAP [33]) that are not covered by any blocks. A new block is then introduced at the center of each point cloud cluster. The other parameters of these new blocks are initialized randomly.
@@
3.2.2 Point-level Optimization
The primitive-based hybrid representation effectively decomposes the shape into parts but performs suboptimally for complex objects. To address this, we introduce a refinement stage for more precise geometry fitting. In this stage, we decouple the constraint between Gaussians and superquadrics, enabling the Gaussians to be freely optimized. Furthermore, to prevent Gaussian components from one block passing through another and disrupting the continuity and plausibility of the part decomposition, we impose a new constraint to minimize the negative signed distance of each Gaussian entering other blocks:
${\mathcal{L}}_{\text{enter}}=\frac{1}{N}\sum_{x\in\Omega}\sum_{m\in\mathcal{M}\setminus\{\delta\}}\text{ReLU}(-D_{m}(x)),$ (14)
where $\Omega$ is the sampled Gaussian set, and $\delta$ represents the identifier of the block to which the Gaussian $x$ belongs. As seen in Fig. 2, the reconstruction becomes more accurate and aligns better with the target shape after this stage.
Figure 3: Qualitative comparison on DTU [17] and ShapeNet [17] . The first two rows are DTU examples, and the last two are ShapeNet examples, respectively. Our method is the only one that provides reasonable 3D part decomposition while capturing detailed geometry.
Figure 4: Qualitative results on BlenderMVS [45] and self-capatured data. We present RGB renderings and decomposed parts from novel views. The top examples are from the BlenderMVS dataset, and the last example is from our captured scenes.
Figure 5: Ablation studies on key strategies. The block-level visual comparisons illustrate the impact of adopting our proposed strategy. The first row shows results without the strategy, and the second with the strategy implemented.
Chamfer distance per scene Mean Mean
Method Input Renderable S24 S37 S40 S55 S63 S65 S69 S83 S97 S105 S106 S110 S114 S118 S122 CD #P
EMS [22] 3D GT 6.32 3.54 2.99 4.30 4.16 4.01 3.75 3.24 4.97 4.34 4.16 7.62 7.58 4.46 4.03 4.65 7.7
MBF [31] 3D GT 3.12 2.66 3.84 2.54 1.59 2.11 2.19 2.01 2.32 2.45 2.17 2.12 3.83 2.02 2.55 2.50 34.1
@@ 3.2.2 Point-level Optimization
Table 1: Quantitative comparison on DTU [17]. The Chamfer distance between the 3D reconstruction and the ground-truth is reported in 15 scenes. The best results are bolded, and the average numbers of primitives found (#P) that are greater than 10 are underlined.
Method Part-aware CD $\downarrow$ PSNR $\uparrow$ SSIM  $\uparrow$ LIPPS $\downarrow$ Time $\downarrow$
[…]
2DGS [16] 0.81 34.07 0.99 0.019 $\sim$ 10 m
PartNeRF [37] 9.59 17.97 0.77 0.246 $\sim$ 8 h
DBW [27] 4.73 16.44 0.75 0.201 $\sim$ 2 h
Ours (Block-level) 4.19 19.84 0.82 0.189 $\sim$ 30 m
Ours (Point-level) 0.95 35.04 0.99 0.015 $\sim$ 40 m
Table 2: Quantitative results on DTU [17]. Our method outperforms all part-aware approaches in image synthesis quality, reconstruction accuracy, and efficiency. Neuralangelo’s results are from the original paper, with all times measured on an RTX 3090 GPU.
@@
Table 3: Quantitative comparison on ShapeNet [6]. We report Chamfer distance and the number of parts. The best results are bolded, and the second-best results are underlined.
@@
Method Part-aware Airplane Table Chair Gun Airplane Table Chair Gun Airplane Table Chair Gun Airplane Table Chair Gun
@@
Ours (Block-level) 2.47 2.15 2.32 1.78 27.94 27.98 24.92 29.95 0.959 0.925 0.906 0.963 0.072 0.129 0.122 0.047
@@
Table 4: Quantitative results on ShapeNet [6]. We report the Chamfer distance and novel view synthesis results across four categories.
4 Experiments
We evaluate our approach on 3D reconstruction quality, view synthesis, shape parsimony, and reconstruction time, comparing it with previous state-of-the-art methods.
4.1 Datasets
We conduct comparative experiments on two public datasets: DTU [17] and ShapeNet [6]. To validate the effectiveness of our method in man-made objects, we construct a ShapeNet subset with four categories: Chair, Table, Gun, and Airplane. The DTU dataset comprises 15 commonly used scenes, and each category in the ShapeNet dataset contains 15 objects. Additionally, to demonstrate our method’s potential with real-world data, more qualitative results are presented on BlenderMVS dataset [45] and self-captured scenes.
4.2 Implementation Details
@@
We build upon the 2DGS [16] architecture and add a non-learnable part attribute to each Gaussian component to enforce the constraint in Eq. 14. In the ShapeNet benchmark, we observe a common degradation in mesh reconstruction as shown in Fig. 7. To address this, we propose Gaussian scale regularization, as validated in Sec. 4.4. The same hyperparameters are used for all experiments. We set the initial number of primitives $M$ to 8. The temperature parameter $\gamma$ in Eq. 9 is 0.005 and the overlapping number $k$ in Eq. 10 is 1.95. We train the hybrid representation for 30k iterations, followed by refinement optimization for another 30k iterations. During block-level optimization, the block-adding operation is carried out at the 5k-th and 10k-th iterations. The loss weights $\lambda_{\text{cov}}$, $\lambda_{\text{over}}$, $\lambda_{\text{par}}$, and $\lambda_{\text{opa}}$ are set to 10, 1, 0.01, and 0.01, respectively. For more details on the dataset, metrics, baselines, and method, please refer to the supplementary materials.
4.3 Evaluations
4.3.1 DTU and Shapenet Benchmark
Geometry Reconstruction. In Tab. 1 and Tab. 3, we compare our geometry reconstruction to SOTA shape decomposition methods on Chamfer distance and training time using DTU and Shapenet dataset. Our method consistently outperforms in all scenes compared to prior works. As shown in Fig. 3, our approach consistently produces interpretable 3D decompositions with further refinement achieving more detailed geometry (the last two columns). MBF [31] achieves a low CD error but at the cost of using significantly more primitives, leading to over-decomposition and the inclusion of meaningless parts. Moreover, as shown in Tab. 2 and Tab. 4, our approach achieves competitive results with advanced Gaussian-based methods(e.g., 2DGS [16]) and surface reconstruction approaches (e.g., Neuralangelo [21]), which are limited to producing unstructured meshes. Notably, our model demonstrates excellent efficiency, with a reconstruction speed approximately 10X faster than PartNeRF and over 3X faster than DBW.
Appearance Reconstruction. Beyond part decomposition, our method enables high-fidelity image synthesis, attributed to integrating Gaussian splatting within superquadric representations. EMS and MBF lack image rendering capabilities, while PartNeRF and DBW yield low-quality view synthesis. In contrast, our method achieves high-quality appearance rendering results, as demonstrated in Tab. 2 and Tab. 4.
Figure 6: Qualitative comparison with SAM-based methods. Our method achieves finer-grained and more distinct part decomposition.